OpenStudy (anonymous):

A university football stadium has 1/6 of its students' seats in the end zones. If tickets are randomly selected and mailed to students, what is the probability that a certain student would get end zone seats at all 5 home games? 5/6 1/30 1/7,776

OpenStudy (anonymous):

Probability that a certain student will get end seat in one match =1/6 Probability that a certain student will get end seat in 5 match =(1/6 )^5= 1/7776 So its (C) :))

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Anytimee (: